OpenStudy (anonymous):

Isaac is moving his herd of 24 sheep to a new pasture.1 acre of land is enough room for 4 of his sheep. How many acres of land does Isaac need to have enough room for all his sheep?

OpenStudy (mathmale):

Have you considered writing a "rate" or "conversion factor?" Multiplying "24 sheep" would produce a result, with units in "acres," that would represent how many acres Isaac would need for his sheep.

OpenStudy (anonymous):

so he'd need 6 acres of land..?

OpenStudy (texaschic101):

1/4 = x/24 4x = 24 x = 6 yep...its 6 acres